please show work! thank you!

(15 points each) Find the general solution: (a) X = = . 3

(b) 1 za = 0 1 1 0 1 1 0 ã. 1 You can use the fact that 13 – 31 – 2 = (1 + 1)(1 – 2)
Answer #1

a ) น Y. hene, A= (13) (2) Now I And] =0 > (PAJ (34) (=6 A2 -4A +10 (A -2 2 2 0 4 = 2,2 t4 where Qų are constant So et We ? aTake M = G = U=-9 so the eigen vector is a Therefore the texty (1) (1) et ta litet, gas ro 9 (1), Gene-toy general selutionFor Araol, AX E-X M2 + 1₂ = -lo ) Mit 2 + 1350 dituz 2-12 Mita2 tloo 4, +42-13 e) 1,4 12 & 11300 Mi = -12-13 Take az a l vec